OpenStudy (anonymous):

the sum of two numbers is 63 and the difference is 9 what are the numbers

OpenStudy (anonymous):

@Hero @perl can one of you help me please I am so confused with this problem

hero (hero):

x + y = 63 x - y = 9

hero (hero):

Solve the system

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

i need the larger number and the smaller number

OpenStudy (anonymous):

@Hero

hero (hero):

Start by adding the equations together.

OpenStudy (anonymous):

how do i add them together?

hero (hero):

By doing this: (x + x) + (y - y) = 63 + 9

OpenStudy (anonymous):

oh ok so what is the next step?

hero (hero):

Combine like terms.

OpenStudy (anonymous):

x^2 +y^2 = 72?

hero (hero):

x + x = 2x y - y = 0

hero (hero):

@RaeBeach9296 are you still here?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so what about the 72?

hero (hero):

You added that correctly.

hero (hero):

So what equation are we left with?

OpenStudy (anonymous):

x+x=2x y+y=0?

hero (hero):

Let's do this a different way: